OpenStudy (anonymous):

Solve for x: 4|2x - 2| + 10 = 34

OpenStudy (solomonzelman):

``` 1. subtract 10 from both sides 2. divide both sides by 4 ``` for now

OpenStudy (solomonzelman):

\(\large\color{slate}{ 4|2x - 2| + 10 = 34 }\) lets start with step #1 (subtracting 10 from both sides) \(\large\color{slate}{ 4|2x - 2| + 10 \color{red}{-10}= 34 \color{red}{-10} }\) what do you then get?

OpenStudy (anonymous):

x = -4, x = 2 ?

OpenStudy (solomonzelman):

the other way around

OpenStudy (solomonzelman):

\(\large\color{slate}{ |2x - 2| =6 }\)

OpenStudy (solomonzelman):

\(\large\color{slate}{ |2(-2) - 2| =|-4-2|=6 }\) \(\large\color{slate}{ |2(4) - 2| =|8-2|=6 }\)

OpenStudy (anonymous):

So x= 4, and x =-2

OpenStudy (adi3):

yes